OpenStudy (pulsified333):

For some positive constant C, a patient's temperature change, T, due to a dose, D, of a drug is given by T=(C/2−D/3)*D^2. What dosage maximizes the temperature change? D= The sensitivity of the body to the drug is defined as dT/dD. What dosage maximizes sensitivity? D=

OpenStudy (pulsified333):

@mathmale @ganeshie8

ganeshie8 (ganeshie8):

So many letters here

ganeshie8 (ganeshie8):

C is a constant T is a function of D

OpenStudy (pulsified333):

Yes

ganeshie8 (ganeshie8):

do you know the standard method to find the mex/min values of some function ?

OpenStudy (pulsified333):

take the derivative and set it to 0?

ganeshie8 (ganeshie8):

Yes, do it

ganeshie8 (ganeshie8):

start by finding the first derivative of T

OpenStudy (pulsified333):

\[\frac{ d(3c-2d) }{ 3 }+d ^{-1/3}\]

OpenStudy (pulsified333):

?

ganeshie8 (ganeshie8):

looks wrong

OpenStudy (pulsified333):

sorry i just had a calc exam 5 hours ago and got 2 hours of sleep last night so im exhausted. But of course, my professor had to assign homework due the day after the test

ganeshie8 (ganeshie8):

its easy, just expand and take the derivaitve : T=(C/2−D/3)*D^2 = (C/2)D^2 - (1/3)D^3 T' = CD - D^2

OpenStudy (pulsified333):

wait the page just refreshed and says its now due Sunday :D. I'll finish this problem with you then hit the sack so i can get a goodnight sleep tonight

ganeshie8 (ganeshie8):

ahh that calc test must have been hard... thats okay :)

OpenStudy (pulsified333):

it was very hard. I think i bombed the test ;(

ganeshie8 (ganeshie8):

haha how sure are you

OpenStudy (pulsified333):

like pretty sure

ganeshie8 (ganeshie8):

does it count too much in the final grade ?

OpenStudy (pulsified333):

not as the midterm and final, i still have a chance at a low b but we will see how the next test goes

ganeshie8 (ganeshie8):

ahh then no big deal, past is past... you can always prepare well for the future tests... cheer up :)

ganeshie8 (ganeshie8):

see if you can finish rest of the problem it should be simple if you know what you're doing..

OpenStudy (pulsified333):

is it \[D=\sqrt{CD}\]

ganeshie8 (ganeshie8):

T' = CD - D^2 right ?

OpenStudy (pulsified333):

yeah

ganeshie8 (ganeshie8):

setting that equal to 0 gives CD - D^2 = 0 D(C - D) = 0 D = 0 or D = C

ganeshie8 (ganeshie8):

zero dosage makes not much sense

OpenStudy (pulsified333):

didn't think about doing it that way

ganeshie8 (ganeshie8):

so D = C must be the maximum

OpenStudy (pulsified333):

okay

ganeshie8 (ganeshie8):

for part b, you need to find the value of D that maximizes the T'

ganeshie8 (ganeshie8):

same thing, differentiate T'

OpenStudy (pulsified333):

you mean take the second derivative?

ganeshie8 (ganeshie8):

T' = CD - D^2 T'' = ?

ganeshie8 (ganeshie8):

T'' is the first derivative of T' and the second derivaitve of T

OpenStudy (pulsified333):

-2D? or -2D+C?

ganeshie8 (ganeshie8):

T' = CD - D^2 T'' = C - 2D

OpenStudy (pulsified333):

ok I thought it was the second one but wasn't sure

OpenStudy (pulsified333):

D=C/2

ganeshie8 (ganeshie8):

looks good!

OpenStudy (pulsified333):

Thanks man, Im gonna head to bed and finish this homework on a fresh head :D

ganeshie8 (ganeshie8):

have good sleep :)